6. Consider an economy that produces and consumes Bread and Automobile. Data for two

different years 2005 and 2010 is given in the following table.

Year 2005 2010Price of Automobiles $ 5000 $ 6000Price of a loaf of bread $10 $20Number of automobiles produced 100 120Number of loaves of bread produced 500,000 400,000



1
Expert's answer
2021-05-16T19:23:00-0400

Nominal GDP (NGDP) = sum of (Current year price x Current year quantity)

Real GDP (RGDP) = sum of (Base year (2005) price x Current year quantity)

GDP Deflator = (NGDP / RGP) x 100

GDP deflator in base year = 100 (since NGDP = RGDP in base year

Inflation rate = % Change in GDP deflator

 

(a)Calculate the nominal and real GDP of 2010.

NGDP, 2005"= 5,000 \times 100 + 10 \times 500,000 = 500,000 + 5,000,000 = 5,500,000"

NGDP, 2010"= 6,000 \times 120 + 20 \times 400,000 = 720,000 + 8,000,000 = 8,720,000"

 

RGDP, 2005"= 5,000 \times100 + 10 \times500,000 = 500,000 + 5,000,000 = 5,500,000"

RGDP, 2010"= 5,000 \times120 + 10 \times 400,000 = 600,000 + 4,000,000 = 4,600,000"

(b) Find the value of GDP Deflator for the year 2010 and interpret.

GDP deflator "= (\frac{8,720,000}{ 4,600,000}) \times 100"

"= 189.57"

Since GDP deflator > 100, this means that the economy experienced inflation between 2005 and 2010.

(c). Calculate the inflation rate in 2010

Inflation rate "= (\frac{189.57}{100}) - 1"

"= 1.8957 - 1"

"= 0.8957= 89.57\%"

(d)compute CPI

cost of basket in 2005 "=5000\times100+10\times500000"

"=500000+5000000=5500000"

cost of basket in 2010 "=6000\times100+20\times500000"

"=600000+10000000=16000000"

CPI in 2005"=(\frac{5500000}{5500000})\times100=100"


CPI in 2010"=(\frac{16000000}{5500000})\times100=290.91"